If logπ‘Ž π‘ž = 1/5 solve the equation , π‘₯ logπ‘Ž π‘ž + (π‘₯ + 2) logπ‘Ž π‘ž2 = 3


1
Expert's answer
2021-11-09T01:03:18-0500

Given: logπ‘ŽΒ π‘ž = 1/5 ...(1)

Now, π‘₯ logπ‘ŽΒ π‘ž + (π‘₯ + 2) logπ‘ŽΒ π‘ž2Β = 3Β 

"\Rightarrow"Β π‘₯ logπ‘ŽΒ π‘ž + (π‘₯ + 2).(2 logπ‘ŽΒ q)Β = 3 ["\because log a^m = m. log a"Β ]

"\Rightarrow"Β Β "\frac{x}{5}+(x+2)\times2\times \frac 1 5=3"Β [From eq. (1)]

"\Rightarrow x+2x+4=15\\
\Rightarrow 3x=11
\\\Rightarrow x =\frac {11}{3}"